OpenStudy (onepieceftw):

How do I factor cosθ out of cos θ - cos θ sin^2θ?

OpenStudy (btaylor):

There's a \(\cos \theta \) term in each side, so you can factor it out by dividing each term by \(\cos \theta\).

OpenStudy (kayders1997):

GCF cos0

OpenStudy (onepieceftw):

There is only one side of this equation.

OpenStudy (btaylor):

sorry, I meant each term of the expression

OpenStudy (kayders1997):

No each term has a cos

OpenStudy (onepieceftw):

So by dividing each term by cosθ I factor it out?

OpenStudy (btaylor):

Yes. \( \cos \theta \left(\cdots \right) \)

OpenStudy (onepieceftw):

So it becomes cosθ(1 - sin^2θ)

OpenStudy (btaylor):

Exactly. Nice work.

OpenStudy (onepieceftw):

Thanks, bud. Appreciate the help!